Lesson 10.2: Goodness of Fit Between Spring 2001 and Spring 2004, the math department at Mt. San Antonio College offered Math wondering if I did it right. Eo came out to 530

Xsquare came out to 163.37

My P Val came out to 0.

Am stuck if I did this right and the conclusion
Answer #1

16) k=7

17)hypothetical proportion=1/7 =0.143

hypothesis are correct:

Applying chi square goodness of fit test:

           relative observed Expected residual Chi square
category frequency Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
L. Advancement Spring 2001 0.143 39 75.71 -4.22 17.803
L. Advancement fall 2001 0.143 37 75.71 -4.45 19.795
27.990 Spring 2002 0.143 112 75.71 4.17 17.390
L. Advancement fall 2002 0.143 67 75.71 -1.00 1.003
27.990 Spring 2003 0.143 159 75.71 9.57 91.614
27.990 fall 2003 0.143 75 75.71 -0.08 0.007
L. Advancement Spring 2004 0.143 41 75.71 -3.99 15.916
total 1.000 530 530 163.528

from above expected frequency Eo =75.71

X2 =163.528

p value =0.0000

as p value is significantly low we reject null hypothesis and conclude that at least one of the proportion is different or enrollment proportions for the 7 semesters are not equal,